84.01g of baking soda, NaHCO3, always contains 22.99 g of sodium, 1.01 g of hydrogen, 12.01 g of carbon, and 48.00 g of oxygen. what percentage of each of these elements is present on baking soda?
1
Expert's answer
2012-10-05T08:14:19-0400
Sodium = (22.99/84.01)*100% = 27.37%

Hydrogen = (1.01/84.01)*100% = 1.20%

Carbon = (12.01/84.01)*100% = 14.30%

Oxygen = (48.00/84.01)*100% = 57.14%